# Testing stimulated emission photon direction

### v3

@@ Abstract
While naively laser only causes excitation of external target, for example STED microscopy demonstrates it can also cause its deexitation, however, under uncommon condition of being prepared as excited. These two causalities are governed by stimulated emission-absorption pair of equations, and swap places in perspective of CPT symmetry, however, it means photon direction of stimulated emission should be opposite to usually assumed. This article discusses various arguments and proposes simple STED-like tests to experimentally determine stimulated emission photon direction(s) once for all, possibly both, to be performed in a near future, already having indirect experimental evidence, like lack of Rabi cycle for free electron laser, or recent ”negative time evidence” experiment in the news. Depending on its result: it should lead to new applications like 2WQC more symmetric quantum computers, or macroscopic violation of CPT symmetry, so far probably tested only in microscopic settings. Also proposed forward/backward/bidirectional photon sources are discussed.
Keywords: lasers, photonics, STED microscopy, CPT symmetry, ring laser, synchrotron radiation, 2WQC
@@ I Introduction
CPT symmetry (of charge conjugation (C) + parity transformation (P) + time reversal (T)) is shown [1] to be necessary for local Lorentz invariant quantum field theories, hence is generally believed to be satisfied by our nature. It has dozens of successful experimental confirmations [2], however, it seems all tests so far were microscopic - come from observations of single particles. Conscious macroscopic test was not found in literature, important especially due to T time symmetry which might lead to new applications - some are proposed and discussed here. For example through system response before impulse, like recently experimentally observed [3] in the news, for example in Scientific American 111 www.scientificamerican.com/article/evidence-of-negative-time-found-….
As in Fig. 1, 2, applying CPT symmetry to scenario ”laser causes target excitation”, it becomes ”CPT(laser) causes CPT(target) deexcitation”, replacing absorption equation with stimulated emission, with photons traveling from CPT(target) to CPT(laser). In contrast, standard textbook view on stimulated emission is energy being kind of knocked by incoming photon - with the opposite direction for the emitted photon. As in Fig. 2, the above consequence of CPT symmetry also allows to conclude the textbook view, hence both can be true.
As there are at least two possibilities for photon direction, the final judge should be experiment. Related tests were earlier proposed using free electron laser [4] (earlier 2009 post), or ring laser [5]. However, they turned out too difficult to organize. This article simplifies test setting to STED-like with just two e.g. diode lasers, fluorescent dye and fast photodetector - determining photon direction through delay. Hence it should be organized soon - to update this article and depending on the outcome: leading to further discussed novel applications, or indicating macroscopic violation of CPT symmetry.
There are also gathered novel arguments, and proposed view on photon sources: using forward/backward/bidirectional beam as in Fig. 3: applying to external target absorption/stimulated emission for unidirectional, or both equations for bidirectional.
Updated article version was motivated by ”negative time evidence” in the news [3], which observed clear signal both before and after the impulse, already indicating absorption-stimulated emission causalities of opposite time delays, already bringing some CPT symmetry suggestions. Also, QInter of our 2WQC group (https://www.qaif.org/2wqc): Vysakh Saji has recently performed the proposed experiment, obtaining unofficial result as $\Delta t\approx 0.0082$ ns for $l\approx d$ setting, this way confirming validity of CPT symmetry in macroscopic scale, to be updated with official results. Second update observes that textbook view can be also concluded from CPT symmetry hence both photon directions can be true, also discusses transparency.
@@
Figure 1: Top: recent widely commented experiment by Steinberg group [3] has observed medium response before and after the applied impulse, proposing ”photon can spend a negative amount of time in an atom cloud” type of explanation, which does not seem physical - in contrast to expected here and looking sufficient: stimulated emission. As impulse used in STED microscopy causes both excitation and deexitation of the target, these two causalities/scenarios swap places in CPT symmetry, hence should have opposite photon direction and sign of delay time as observed. Bottom: proposed simplified setting/test, recreating STED situation, but focused on temporal resolution. Continuous laser excites fluorescent dye being approximately 3-level medium, and impulse laser causes its deexcitation through stimulated emission - the question to test is their photon direction, which can be found from $\Delta t=(d\pm l)/c$ delay between laser impulse and observed reduced intensity by fast photodetector focused on dye spontaneous emission. Textbooks assumption would mean ${}^{\prime}+^{\prime}$ sign here, while CPT symmetry requires ${}^{\prime}-^{\prime}$ sign instead, hence this is also macroscopic test of CPT symmetry, complementing dozes of earlier microscopic tests [2].
@@ II STED microscopy based tests
Stimulated emission depletion (STED) microscopy [6] improves resolution by reducing size of excited dot with donought-like depletion beam from second laser as in Fig. 3. It uses fluorescent dye being approximately 3 level medium: excited with one laser from 0 to 2, then fast relaxation to 1, then deexcitation from 1 to 0 with slightly lower wavelength - usually through just spontaneous emission observed by microscope lens.
In STED we additionally speedup this deexcitation through stimulated emission by second: depletion laser of frequency for $1\to 0$ transition - reducing the number of photons from this region monitored by microscope lens focused on spontaneous emission, this way allowing to increase microscope resolution.
While spontaneous emission photons should have a random direction, the question of concern is direction of such stimulated emission photons: toward ()($\leftarrow$)( ← ) or outward ()($\rightarrow$)( → ) the laser? Textbooks assume the latter, discussion here based on CPT symmetry suggests the former. There can be also both, or a different angle, what also could be tested experimentally.
@@ II-A Delay STED-like test
Assuming there are only these two options: ()($\leftarrow$)( ← ) or ()($\rightarrow$)( → ), the simplest way to distinguish them seems delay. There is literature discussing delay in STED microscopy, like [7, 8, 9], however, data they provide seem insufficient to uniquely answer this direction question.
@@
Hence photodetector should be connected to oscilloscope, triggered by depletion laser impulse. Its observed delay $\Delta t$ contains positive contribution $d/c$ for light path from dye to detection, and ±$\pm l/c$ contribution from laser-dye path, of sign for one of the two assumed possibilities - which could be determined from this delay, e.g. from oscilloscope snapshots planned to be included in the next version of this article.
@@ II-B Alternative STED-like tests
[…]
As discussed further, additional tests could be using unidirectional e.g. ring laser: its forward beam should not allow for depletion here, its backward beam should allow - what could be also tested in the future.
@@
Rabi cycle could be used instead, by applying impulse laser and monitoring/analyzing situation at impulse boundaries.
@@ III CPT symmetry in photonics arguments
Let us gather here some arguments for the expected ()($\leftarrow$)( ← ) possibility mainly based on CPT symmetry. However, as in Fig. 2, it seems to also imply textbook ()($\rightarrow$)( → ) case, hence both can be true - practically this is test of existence of ()($\leftarrow$)( ← ) effect.
@@ III-A Bidirectional and forward/backward beams
Looking at such unidirectional ring laser from CPT perspective, photon trajectory would be reversed, causing excitation of different (backward) target - increasing $N_{2}$ toward minus time, hence decreasing $N_{2}$ in standard time perspective - as in stimulated emission. Therefore, this CPT symmetry switches the two equations - in contrast e.g. to diode lasers in STED microscopy, seems only one equation at the time acts on the forward/backward targets - allowing to split beams them into intuitively forward/backward beam acting with separated absorption/stimulated emission equation alone, combined into bidirectional beam in standard e.g. diode laser.
@@
As noticed in Fig. 2, forward/backward beams might have completely different transparency as they act only on $N_{1}$ deexcited / $N_{2}$ excited atoms. As everyday matter usually has $N_{2}/N\approx 0$, backward beam might have much better penetration, allowing for novel applications.
@@
Figure 3: Top: ring laser uses optical isolator to enforce unidirectional photon circulation, which would be reversed in CPT perspective - switching the action of absorption and stimulated emission equations, suggesting only one of them should act on external targets - let us refer to them as forward/backward beam. Analogously for synchrotron radiation, wiggler, undulator, free electron laser - reversing electron and photon trajectory in CPT perspective. Center: STED microscopy (image source) showing that, at least for diode lasers, both equations act on external target, hence we can call it bidirectional beam. We usually focus on absorption by external target, because stimulated emission requires its initial excitation, while without such preparation usually $N_{2}/N\approx 0$. In Rabi cycle laser periodically causes excitation and deexitation of target, suggesting requirement of bidirectional beam. Its experimental support is lack of Rabi oscillations for free electron laser [13] - leading to $N_{2}\approx N$ nearly complete excitation in Autler-Townes effect (dynamic Stark), instead of $N_{2}\approx N/2$ for damped Rabi cycle. Bottom: while CPT symmetry of physics is required by local Lorentz-invariant QFT, e.g. switching forward/backward beams, in theory general relativity allows its realization by travelling through nonorientable wormhole [12], bringing tangible thought experiment valuable for intuitions.
@@ III-B Experimental consequences like Rabi cycle
To test the above with STED-like setting, as mentioned we can replace depletion laser with forward/backward beam - only backward should lead to depletion. If stimulated emission-absorption use opposite photon trajectories, we could also try to split them exploiting delay with fast shutter or switch, or maybe exploiting transparency absorbing only forward beam.
Another type of consequence is $N_{1}-N_{2}$ population in external target of laser. For bidirectional acting with both equations, we could at most reach 1/21/$2-1/$ 2 - 1 / 2 population, and it is indeed well known as the final population of damped Rabi cycle [14]: using periodically stimulated emission and absorption equations. In Fig. 4 there is suggested hydrodynamic analog of Rabi cycle (another can be found in [15]), which should exist through EM-hydrodynamics analogy as written, like the shown resonator coupled with external object exchanging momentum in both directions, what might be also appropriate intuition for bidirectional lasers, using cyclically positive and negative radiation pressure [16]. Rabi cycle in CPT perspective should be practically the same - bringing another argument for photons travelling there only between laser and target.
However, using only forward beam, we should be able to reach $N_{2}\approx N$ nearly complete excitation, and such effect has turned out to be known in literature: despite many attempts and theoretical prediction, they were not able to obtain Rabi oscillation with free electron laser [13] - instead getting Autler-Townes effect (dynamics Stark) leading to $N_{2}\approx N$, instead of $N_{2}\approx N/2$ for damped Rabi oscillations.
@@ III-C Some possible new applications
If confirmed, the above considerations could lead to many new applications, especially of currently neglected backward beam, like:
@@
  • If we could send information this way, there could be e.g. built time-loop computers (Section V of [4]) for example using backward beam to send chip output back to its original input with negative time delay, also in theory allowing to solve NP problems. Connecting multiple such negative time transmitters e.g. into a cycle, could allow to also reach macroscopic negative time differences.
  • Backward camera e.g. exploiting transparency or for astronomical observations - focused on target’s backward beam, while standard camera is focused on its forward beam. It would replace CMOS/CCD matrix behind the lens, with matrix of continuously excited e.g. fluorescent dye, plus photodetectors monitoring their spontaneous emission.
  • Backward beam should be transparent for $N_{2}=0$ medium, what might allow for much better penetration than forward (transparent for $N_{1}=0$) - for example for backward camera or communication through obstacles.
  • Backward beam for attacks on quantum cryptographic protocols like BB84 [18] - use external laser to additionally stimulate emission inside communication laser to steal/intercept such additional photons, which could be read e.g. by passing through control of CNOT gate - affecting second qubit to be measured.
  • Backward beam for observations actively increasing deexcitation rate of excited objects, e.g. for thermal imaging,
  • Backward beam for 3D printing/reversed lithography: excite material and precisely stimulate its deexcitation - controlling position and time, also frequency to choose one of multiple possible transitions, e.g. some high energy transition to locally reduce temperature.
  • Synchrotron sources usually have much higher photon energy, its backward beam could e.g. increase probability of selected nuclear decays/transitions: producing photons of chosen energy. Maybe also for other particle events: which CPT transform use such photons coming from e.g. synchrotron.

### v4

@@ Abstract
While naively laser only causes excitation of external target, for example STED microscopy or backward ASE in optical amplifiers, demonstrate it can also cause its deexitation, however, under uncommon condition of being prepared as excited. These two causalities are governed by absorption-stimulated emission pair of equations, and swap places in perspective of CPT symmetry, however, it means photon direction of stimulated emission should be opposite to usually assumed. This article discusses various arguments and proposes simple direct tests to experimentally verify existence of such backward photon trajectories, complementing consequent forward textbook trajectories. There are already indirect experimental arguments, like backward ASE, lack of Rabi cycle for free electron laser, or recent ”negative time evidence” experiment in the news. Depending on the result: it should lead to new applications like 2WQC more symmetric quantum computers, or macroscopic violation of CPT symmetry, so far probably tested only in microscopic settings. Also proposed forward/backward/bidirectional photon sources are discussed.
Keywords: lasers, photonics, STED microscopy, backward ASE, CPT symmetry, ring laser, synchrotron radiation, 2WQC
@@ I Introduction
Figure 1: Top: recent widely commented experiment by Steinberg group [1] has observed medium response (measured phase of 85Rb) before and after the applied impulse, proposing ”photon can spend a negative amount of time in an atom cloud” type of explanation, which does not seem physical. In contrast, as discussed here, this might be a result of just stimulated emission, required as some of the atoms were excited by magneto-optical trap (MOT) lasers, and if being CPT analog of absorption, should have opposite delay sign.
@@
CPT symmetry (of charge conjugation (C) + parity transformation (P) + time reversal (T)) is shown [2] to be necessary for local Lorentz invariant quantum field theories, hence is generally believed to be satisfied by our nature. It has dozens of successful experimental confirmations [3], however, it seems all tests so far were microscopic - come from observations of single particles. Conscious macroscopic test was not found in literature, important especially due to T time symmetry which might lead to new applications - some are proposed and discussed here. For example through system response before impulse, like recently experimentally observed by Steinberg group in the news [1]111 www.scientificamerican.com/article/evidence-of-negative-time-found-….
@@
Figure 2: Top: CPT symmetry says that if there exists ”laser causes excitation” scenario, also ”laser causes deexciation” should exist. Being CPT symmetry analogs, they should have opposite photon direction and delay sign - we would like to verify it experimentally and apply if successful. For example in STED microscopy diode lasers cause both. Bottom: proposed STED-like test of this photon direction(s): continuous laser excites fluorescent dye being approximately 3-level medium, and impulse laser causes its deexcitation through stimulated emission - its photon direction can be found from $\Delta t=(d\pm l)/c$ delay between laser impulse and observed reduced intensity by fast photodetector focused on dye spontaneous emission. Textbooks assumption would mean ${}^{\prime}+^{\prime}$ sign here, while CPT symmetry requires ${}^{\prime}-^{\prime}$ sign instead, hence this is also macroscopic test of CPT symmetry, complementing dozes of earlier microscopic tests [3].
Figure 3: In commercially available e.g. erbium-doped optical fiber amplifiers (EDFA or alternatives) [4] there is observed strong backward ASE (amplified spontaneous emission also called superluminescence) - increased emission from pumped doped fiber toward laser, exactly as in the sought CPT analog of ”laser causes excitation” effect. Analogously to STED-like test, we can verify (and apply if successful) this photon direction by measuring delay between the impulse and reduced population of excited atoms of external target, e.g. by monitoring its spontaneous emission. For this purpose, the shown standard forward optical isolator should be removed, or better rotated to backward position. As erbium has very long millisecond-scale lifetimes (should be shortened for practical applications), there might be required cascade photodiode.
As in Fig. 2, applying CPT symmetry to ”laser causes target excitation” scenario, it becomes ”CPT(laser) causes CPT(target) deexcitation”, replacing absorption equation with stimulated emission, having photons traveling from CPT(target) to CPT(laser), also reversing time delay, what might explain observations of negative time delays, also e.g. [5], finally being a consequence of large number of single emissions/absorptions.
We would like to verify existence of such CPT analog of ”laser causes excitation” scenario with backward photon trajectory, for example in STED-like setting as in Fig. 2. Well known backward ASE in common optical amplifiers, as in Fig. 3, seems exactly the sought scenario, also providing practical realization of such test, and maybe also discussed further potential applications thanks to wide commercial availability of optical fiber hardware.
In contrast, as in Fig. 4, standard textbook view on stimulated emission is energy being kind of knocked forward by incoming photon - with the opposite direction for the emitted photon. There are lots of experimental evidence for existence of such process. While it seems in disagreement with CPT symmetry, bottom of Fig. 4 shows it can be viewed as a consequence of the sought: CPT analog of ”laser causes excitation” scenario.
As there are at least these two possibilities for stimulated emission photon directions, the final judge should be experiment. Related tests were earlier proposed using free electron laser [6] (earlier 2009 post), or ring laser [7]. However, they turned out too difficult to organize. This article simplifies test setting to STED-like, and later to optical amplifier - determining photon direction through delay. Hence it should be organized soon - to update this article, and depending on the outcome: leading to further discussed novel applications, or indicating macroscopic violation of CPT symmetry.
Absorption/stimulated emission equations are switched in CPT perspective, and e.g. diode laser in STED microscope acts with both on external target (bidirectional beam), also Rabi cycle requires both. There is a natural question if we can separate them - there will be shown arguments that e.g. optical isolator or synchrotron radiation allows to split them - into forward/backward unidirectional beam as in Fig. 6.
Multiple updates of this article were mainly motivated by finding additional indirect experimental evidence in the literature (e.g. for Fig. 1, 3, 5), transparency with textbook view as consequent, convenient optical amplifier realization, and adding multiple potential applications. Also, QInter of our 2WQC group (https://www.qaif.org/2wqc): Vysakh Saji has recently performed the proposed experiment, obtaining unofficial result as $\Delta t\approx 0.0082$ ns for $l\approx d$ setting, this way confirming validity of CPT symmetry in macroscopic scale, to be updated with official results.
[…]
Figure 5: Other type of argument for time/CPT symmetry of lasers are pre-pulses generated by post-pulses, observed (e.g. [8, 9]) for extremely high energy lasers based on chirped-pulse amplification (CPA) going through nonlinear medium (being first stretched then compressed). Theoretical analysis is based on Fourier transform in time, being symmetrized by nonlinearities. However, decomposing it into individual absorptions/emissions, negative delays should again come from being CPT analogs, hence having opposite delay signs.
@@ II STED microscopy based tests
Stimulated emission depletion (STED) microscopy [10] improves resolution by reducing size of excited dot with donought-like depletion beam from second laser as in Fig. 6. It uses fluorescent dye being approximately 3 level medium: excited with one laser from 0 to 2, then fast relaxation to 1, then deexcitation from 1 to 0 with slightly longer wavelength - usually through just spontaneous emission observed by microscope lens.
In STED we additionally speedup this deexcitation through stimulated emission by second: depletion laser of frequency for $1\to 0$ transition - actively reducing the number of photons from doughnut-shaped region. This way the final excited dot is smaller, increasing microscope resolution for lens focused on stimulated emission of the dye.
While spontaneous emission photons should have a random direction, the question of concern is direction of stimulated emission photons: ()($\rightarrow$)( → ) forward ano/or ()($\leftarrow$)( ← ) backward? Textbooks assume forward, discussion here based on CPT symmetry suggests backward as fundamental, and forward as its consequence (Fig. 4). Therefore, both can be true, the question to verify is existence of backward ()($\leftarrow$)( ← ).
@@ II-A Delay STED-like test
Assuming there are only these two options: ()($\leftarrow$)( ← ) or ()($\rightarrow$)( → ) and we would like to test existence of the former, the simplest way to distinguish them seems delay. There is literature discussing delay in STED microscopy, like [11, 12, 13], however, data they provide seems insufficient to uniquely answer this question.
@@
Hence (fast) photodetector should be connected to oscilloscope, triggered by depletion laser impulse. Its observed delay $\Delta t$ contains positive contribution $d/c$ for light path from dye to detector, and ±$\pm l/c$ contribution from impulse laser-dye path, of sign for one of the two assumed possibilities - which could be determined from this delay, e.g. from oscilloscope snapshots planned to be included in the next version of this article.
@@
As in Fig. 3, standard commercially available EDFA fiber optics equipment seems very convenient for analogous test and further applications if successful (suggested by Michał Markiewicz). The sought target $\to$ laser photon trajectories caused by laser seems exactly the backward ASE effect, well known in literature - usually unwanted and removed by optical isolator. We can test these photon trajectories by removing or better reversing this optical isolator, adding (rather fast) photodetector filtered to observed the stimulated emission, and observing delay of its reduction as response for impulse of deexcitation laser. In comparison to STED setting, $d/c$ propagation time is between EDF and detector, $l/c$ between EDF and impulse laser - easily extended by delay line.
@@ II-B Alternative STED-like tests
As discussed further, additional tests could use unidirectional e.g. ring laser (or standard with added optical isolator): its forward beam should not allow for depletion here, its backward beam should allow - what could be also tested in the future.
@@
Rabi cycle could be used instead, by applying impulse laser and monitoring/analyzing situation at impulse boundaries, or including optical isolator on the way.
@@ III CPT symmetry in photonics arguments
Let us gather here some arguments for the expected ()($\leftarrow$)( ← ) possibility mainly based on CPT symmetry. However, as in Fig. 4, it seems to also imply textbook ()($\rightarrow$)( → ) case, hence both can be true - hence practically this is a test of existence of ()($\leftarrow$)( ← ) effect.
@@ III-A Bidirectional and forward/backward beams
Figure 6: Top: ring laser uses optical isolator to enforce unidirectional photon circulation, which would be reversed in CPT perspective - switching the action of absorption and stimulated emission equations, suggesting only one of them should act on external targets - let us refer to them as forward/backward beam. Analogously for just adding optical isolator, or synchrotron sources like wiggler, undulator, free electron laser - reversing electron and photon trajectory in CPT perspective. Center: STED microscopy (image source) showing that, at least for diode lasers, both equations act on external target, hence we can call it bidirectional beam. We usually focus on absorption by external target, because stimulated emission requires its initial excitation, while without such preparation usually $N_{2}/N\approx 0$. In Rabi cycle laser periodically causes excitation and deexitation of target, suggesting requirement of bidirectional beam. Its experimental support is lack of Rabi oscillations for free electron laser [14] - leading to $N_{2}\approx N$ nearly complete excitation in Autler-Townes effect (dynamic Stark), instead of $N_{2}\approx N/2$ for damped Rabi cycle. Bottom: while CPT symmetry of physics is required by local Lorentz-invariant QFT, e.g. switching forward/backward beams, in theory general relativity allows its realization by travelling through nonorientable wormhole [15], bringing tangible thought experiment valuable for intuitions.
@@
Looking at such unidirectional ring laser from CPT perspective, photon trajectory would be reversed, causing excitation of different (backward) target - increasing $N_{2}$ toward minus time, hence decreasing $N_{2}$ in standard time perspective - as in stimulated emission. Therefore, this CPT symmetry switches the two equations - in contrast e.g. to diode lasers in STED microscopy, seems only one equation at the time acts on the forward/backward targets - allowing to split beams into intuitively forward/backward beam acting with separated absorption/stimulated emission equation alone, combined into bidirectional beam in standard e.g. diode laser.
@@
As noticed in Fig. 4, forward/backward beams might have completely different transparency as they act only on $N_{1}$ deexcited / $N_{2}$ excited atoms. In other words, while textbook view on stimulated emission means amplification of forward beam by going through excited medium, its CPT analog is amplification of backward beam by going through deexcited medium (as in this Figure). Therefore, as everyday matter usually has $N_{2}/N\approx 0$, backward beam might have much better penetration, allowing for novel applications.
@@ III-B Experimental consequences like Rabi cycle
To test the above with STED-like setting, as mentioned we can replace depletion laser with forward/backward beam e.g. by optical isolator - only backward should lead to depletion. If stimulated emission-absorption use opposite photon trajectories, we could also try to split them exploiting delay with fast shutter or switch, or maybe exploiting transparency absorbing only forward beam.
Another type of consequence is $N_{1},N_{2}$ population enforced by laser in its external target. For (symmetric) bidirectional beam acting with both equations, we could at most reach 1/21/$2-1/$ 2 - 1 / 2 population, and it is indeed well known as the final population of damped Rabi cycle [18]: using periodically stimulated emission and absorption equations. In Fig. 7 there is suggested hydrodynamic analog of Rabi cycle (another can be found in [19]), which should exist through EM-hydrodynamics analogy as written, like the shown resonator coupled with external object exchanging momentum in both directions, what might be also appropriate intuition for bidirectional lasers, using cyclically positive and negative radiation pressure [20]. Rabi cycle in CPT perspective should be practically the same - bringing another argument for photons travelling there only between laser and target.
However, using only forward beam, we should be able to reach $N_{2}\approx N$ nearly complete excitation, and such effect has turned out to be known in literature: despite many attempts and theoretical prediction, they were not able to obtain Rabi oscillation with free electron laser [14] - instead getting Autler-Townes effect (dynamics Stark) leading to $N_{2}\approx N$, instead of $N_{2}\approx N/2$ for damped Rabi oscillations. There should be also possible asymmetric bidirectional beams e.g. by adding multiple beams, or applying imperfect optical isolators - with different powers of forward/backward beams, for example leading to damped Rabi cycle approaching a different population.
@@ III-C Some possible new applications
If confirmed, the above considerations could lead to many new applications of e.g. standard optical fiber equipment, especially for the currently neglected backward beam, like:
@@
  • If we could send information this way, there could be e.g. built time-loop computers (Section V of [6]) for example using backward beam to send chip output back to its original input with negative time delay, also in theory allowing to solve NP problems. Connecting such multiple negative delay transmitters e.g. into a cycle, could allow to also reach macroscopic negative time differences.
  • Backward camera e.g. exploiting transparency or for astronomical observations - focused on target’s backward beam, while standard camera is focused on its forward beam. It would replace CMOS/CCD matrix behind the lens, with matrix/surface of continuously excited approximately 3-level medium like dye, plus some monitoring of its population map - e.g. through spontaneous emission, or EIT (electromagnetically induced transparency).
  • Backward beam might be transparent for $N_{2}=0$ medium, what could allow for much better penetration than forward (transparent for $N_{1}=0$) - e.g. for backward camera, various scanners, communication through obstacles.
  • Backward beam for attacks on quantum cryptographic protocols like BB84 [23] - use external laser to additionally stimulate emission inside communication laser (as backward ASE) to steal/intercept such additional photons, which could be read e.g. by passing through control qubit of CNOT gate - affecting the second qubit to be measured.
  • Backward beam for observations actively increasing deexcitation rate of excited objects, e.g. for thermal imaging.
  • Backward beam to stimulate specific chemical processes: releasing photons of chosen energy/polarization, e.g. for chemical/technological processes, energy production, or medical therapies - e.g. locally reduce lifetime of ATP molecule (into thermal energy also heating), or inhibit biochemical processes by causing deexcitation of crucial molecules, for example with multiple beams intersecting in cancer tissue, especially if better transparency is confirmed.
  • Backward beam for 3D printing/reversed lithography: excite material and precisely stimulate its deexcitation - controlling position and time, also frequency to choose one of multiple possible transitions e.g. some high energy transition to locally reduce temperature, maybe polarization.
  • Synchrotron sources usually have much higher photon energy, its backward beam could e.g. increase probability of selected nuclear decays/transitions: producing photons of chosen energy, polarization. Maybe also for other particle events: which CPT transform requires such photons incoming from e.g. synchrotron.